Shift report 20010908 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Sep 08 2001 - 00:16:29 EDT

  • Next message: Apache: "Shift report 20010908 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Sep  8 00:16:29 2001, Shift: 16:00-24:00
    Supervisor: K. Hagel, with: 
    Log book#8 pages 91 - 93
    
    Shift Summary:
    ------------------------------------------------------------
    5 minutes of beam during the entire shift
    Then: Security from PASS division A lost.  RHIC zones
    6z1 and 6z2 need to be reswept
    
    At shift end, monitor reads that RHIC has been resecured;
    Preparing from 56 bunch ramp.
    
    Did a long DVM calibration run to try to get statistics
    for TPM1 to verify drift velocity.  That was run 4635.
    We can see if it was enough.  Plots with centroids notated
    are pasted in logbook.
    
    NOTE: During shift check, I noticed that Ar working is
          empty (reserve bottles are OK) and that the N2 working
          is nearly empty (reserve bottles also OK)  I don't
          think this is urgent, but could become so if not
          addressed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Sep 08 2001 - 00:17:06 EDT